OpenStudy (isbella):

1)integral ycos(x^2) dx i dont know how to do this one can someone help me out

OpenStudy (isbella):

\[\int\limits ycos(x^2) dx\]

geerky42 (geerky42):

And what is y equal to?

OpenStudy (isbella):

this is part of the question on my hw sheet of double integarl \[\int\limits_{0}^{2}\int\limits_{y^2}^{4} ycos(x^2) dxdy\]

geerky42 (geerky42):

I'm new to double integral so I could be wrong. but I think I know how to do it. First integral is in respect of x so you just treat y as constant for first integral. \[\int\limits_{0}^{2} \int\limits_{y^2}^{4}y\cos(x^2)dxdy = \int\limits_{0}^{2}y \left[ \int\limits_{y^2}^{4} \cos(x^2)dx\right]dy\] Do you know what this is?\[\int\cos(x^2)dx \]

OpenStudy (isbella):

nope

ganeshie8 (ganeshie8):

switch the order to dydx

ganeshie8 (ganeshie8):

\(\mathbb{\int\limits_{0}^{2}\int\limits_{y^2}^{4} ycos(x^2) dxdy = \int\limits_{0}^{4}\left(\int\limits_{0}^{\sqrt{x}} ycos(x^2) dy\right)dx}\)

OpenStudy (ipwnbunnies):

Yes, this is a good problem in which switching the order of integration is cha-ching. ;)

OpenStudy (isbella):

can some one please show me the steps i would be very greatful